# CLEAR: Calibrated Learning for Epistemic and Aleatoric Risk

### v1

@@ Appendix A Related Literature
In the main paper, Section 1 provides a high-level overview of uncertainty quantification in ML. Section 2.2 contains a quick overview of epistemic uncertainty, and Section 2.3 discusses the aleatoric uncertainty. In this appendix, we discuss these concepts in more depth. For an introduction to epistemic and aleatoric uncertainty, see [11], which provides a comprehensive overview.
A.1 Literature that implicitly assumes $\gamma_{1}=1$ when combining epistemic and aleatoric uncertainty
@@
When comparing against UACQR, we see in Tables 18, 19 and 20 that the performance of CLEAR is much more reliable across the considered datasets and metrics. E.g., on the airfoil, energy efficiency and naval propulsion CLEAR is approximately 2 to 3 times better in metrics such as Quantile Loss, NCIW and Average Interval Score Los, while UACQR never outperforms CLEAR by more than 27% in any of the datasets in any of the considered metrics. Sometimes UACQR-P can also output infinitely wide predictive intervals [19, p. 5], which we observed for one dataset in our experiments. CLEAR performs better than both versions of UACQR in 14 out of 17 datasets. These large difference in performance in Section E.3 can partially be explained by our approach to fit the aleatoric uncertainty on the residuals. However, in Sections E.1 and E.2 every method uses exactly the same module for estimating the point prediction $\hat{f}$, the uncalibrated epistemic uncertain $q^{\text{epi}}$ and the uncalibrated $q^{\text{ale}}$. This allows us to only compare the high-level concepts of fixing $\gamma_{1}=1$ versus optimizing both $\gamma_{1}$ and $\lambda$. In this case the gap narrows down substantially, but still CLEAR is more robust. Fixing $\gamma_{1}=1$ can never outperform CLEAR by more than 10% on any dataset with respect to any metric. We hypothesize that CLEAR is more stable and robust since it can compensate for shortcomings of the base models more easily: If the aleatoric uncertainty is overestimated or underestimated, CLEAR can correct its scale by adjusting $\gamma_{1}$.
A.2 Literature that implicitly assumes $\lambda=1$ when combining epistemic and aleatoric uncertainty
Most of the literature that combines epistemic and aleatoric uncertainty implicitly assumes that $\lambda=1$, when they simply combine epistemic and aleatoric uncertainty in the ration 1:1. On top of the resulting uncertainty once can use (conformal) calibration, which corresponds to CLEAR’s calibration of $\gamma_{1}$. However, in contrast to CLEAR, they do not rebalance the ratio of epistemic and aleatoric uncertainty with $\lambda$.
In practice, it is possible to underestimate the aleatoric uncertainty and overestimate epistemic uncertainty. This results in a one-dimensional calibration via $\gamma_{1}$, and consequently, narrow intervals in regions of dominating aleatoric uncertainty or in wide intervals in regions of dominating epistemic uncertainty. $\gamma_{1}$ alone cannot solve this alone. CLEAR can deal well with such situations by compensating for such an imbalance via $\lambda$.
@@ A.2.1 Deep Ensembles
While it is quite common to refer to deep ensembles (DE), whenever one uses an ensemble of neural networks (NNs) for uncertainty estimation, it is important to note that [22] introduced DE as a method which both estimates epistemic and aleatoric uncertainty. For regression, they train each NN with 2 outputs estimating $\mu$ and $\sigma$ via a Gaussian Maximum-Likelihood-loss (as in [35]), where $\sigma$ is responsible for the aleatoric uncertainty, which they refer to as “ambiguity in targets y for a given x”. Moreover, they use the ensemble diversity to estimate epistemic uncertainty, which they refer to as “model uncertainty”. Although DE is mainly known for its ensembling approach, [22, Table 2 in Appendix A.1] clearly shows that both the aleatoric and epistemic parts are crucial in terms of empirical performance. In their paper, the authors do not apply any calibration on top, i.e., as it is implicitly assumed that $\gamma_{2}=\gamma_{1}=1=\lambda$. However, it is common to apply (conformal) calibration on top. The commonly used calibration techniques only calibrate $\gamma_{1}$, while keeping $\lambda=1$ fixed, in contrast to CLEAR. In particular, for DE, both the diversity of the ensemble and the bias on aleatoric are very sensitive to various hyperparameters. One type of uncertainty may be strongly under-estimated. In contrast, the other type is strongly over-estimated, which motivates the need to calibrate explicitly calibrate $\lambda$ in a data-driven way.
@@
The original paper [22] is written as if applying adversarial attacks is an integral part of DEs and one of the paper’s main contributions. However, to the best of our knowledge, many practitioners refer to “DEs” without implying adversarial attacks during training, and adversarial attacks during training are seen as an optional add-on to DEs but not an important part of DEs. Furthermore, the regression results in [22, Table 2 in Appendix A.1] do not suggest that adversarial attacks during training are particularly beneficial to the performance.
@@
Intuitively, one should estimate high epistemic uncertainty if there is a significant disagreement among the ensemble’s predictions and small epistemic uncertainty if they agree. While [24, Chapter 13] and [23] suggest using quantiles, [22] suggest using the empirical standard deviation to estimate DE’s disagreement among ensemble’s predictions. It seems plausible that in [22]’s case of 5 ensemble members, using the standard deviation and some Gaussian assumptions can be more appropriate while in [24, Chapter 13]’s and [23]’s case of 100 or more ensemble members, quantiles can be more accurate.
@@
There exist several modifications of DE that, for example, promote the ensemble’s diversity on the function space via an additional loss term during training [76], ensemble over multiple different hyperparameters [77], or reduce the computational training cost [17, 27, 78, 79, 19, 23].
@@ A.2.2 Monte Carlo Dropout
[27] originally studied Monte Carlo Dropout (MC Dropout) without explicitly modeling aleatoric uncertainty, which was then extended by [17] to also explicitly model aleatoric paper. While [27] see MC Dropout as an approximation of a Bayesian neural network, one can also see it as another ensemble method. The main difference to DE is that MC Dropout only trains one NN with dropout and obtains an ensemble after training by randomly setting weights of the model to zero at inference time. Analogously to DE, MC Dropout also adds epistemic and aleatoric uncertainty in the ratio 1:1 with the same disadvantages as described before in Section A.2.1.
@@
[80] empirically concludes that ensembles have the best uncertainty and disentangling behavior of epistemic and aleatoric uncertainty. In their paper, the authors do not use any form of calibration. This would correspond to $\gamma_{2}=\gamma_{1}=1=\lambda$ in our notation. Their paper suggests a different loss function, which they call $\beta$-NLL, for training to mitigate the underestimation of aleatoric uncertainty to some extent in their experimental setting without comparing this approach to calibration. In [80, Figure 6], one can observe that even without the $\beta$-NLL, both epistemic and aleatoric uncertainty already have a good shape (that is, good relative uncertainty, see Appendix H) for deep ensembles (DE). The main problem of DE in [80, Figure 6] is that the epistemic uncertainty is too small by a very large factor (that is, poor absolute scale of uncertainty, see Appendix H), while aleatoric uncertainty has already almost the correct scaling. From our perspective, applying CLEAR on their DE would probably largely fix their problem of DE if CLEAR chooses $\gamma_{1}\approx 1$ and $\lambda\gg 1$. However, they show that for this specific experiment, $\beta$-NLL also fixes the problem. In general, where one suspects that the aleatoric or the epistemic uncertainty might be too small or too large across the domain, we strongly recommend simply applying CLEAR on top of the already trained ensemble instead of retraining all the models with a new training pipeline. The concept of CLEAR can be implemented in a few minutes and calibrates an already trained ensemble in a few seconds. We considered an interesting open problem to study if $\beta$-NLL can improve the relative epistemic and aleatoric uncertainty (see Appendix H).
@@
[14] also combines epistemic and aleatoric uncertainty in the ratio 1:1 and applies a single constant (corresponding to $\gamma_{1}$ in our notation) to scale the total predictive uncertainty, which corresponds to fixing $\lambda=1$ resulting in the same potential for problems as mentioned before.
@@ A.3.1 Gaussian Process Regression (GPR)
A.4 Bayesian Neural Networks (BNNs)
@@ A.4.2 TabPFN
A.5 Conformal Literature that does not account for Epistemic and Aleatoric Uncertainty

### v2

@@ Appendix A Related Literature
In the main paper, Section 1 provides a high-level overview of uncertainty quantification in ML. Section 2.2 contains an overview of epistemic uncertainty, and Section 2.3 discusses the aleatoric uncertainty. In this appendix, we discuss these concepts in more depth. For an introduction to epistemic and aleatoric uncertainty, see [11], which provides a comprehensive overview.
A.1 Literature that implicitly assumes $\gamma_{1}=1$
@@
When comparing against UACQR, we see in Tables 24, 25 and 26 that the performance of CLEAR is much more reliable across the considered datasets and metrics. For example, on the airfoil, energy efficiency and naval propulsion CLEAR is approximately 2 to 3 times better in metrics such as Quantile Loss, NCIW and Average Interval Score Loss, while UACQR never outperforms CLEAR by more than 27% in any of the datasets in any of the considered metrics. Sometimes UACQR-P can also output infinitely wide predictive intervals [19, p. 5], which we observed for one dataset in our experiments. CLEAR performs better than both versions of UACQR in 14 out of 17 datasets. These large difference in performance in Section F.3 can partially be explained by our approach to fit the aleatoric uncertainty on the residuals. However, in Sections F.1 and F.2 every method uses exactly the same module for estimating the point prediction $\hat{f}$, the uncalibrated epistemic uncertainty $q^{\text{epi}}$, and the uncalibrated $q^{\text{ale}}$. This allows us to only compare the high-level concepts of fixing $\gamma_{1}=1$ versus optimizing both $\gamma_{1}$ and $\lambda$. In this case the gap narrows down substantially, but still CLEAR is more robust. Fixing $\gamma_{1}=1$ can never outperform CLEAR by more than 10% on any dataset with respect to any metric. We hypothesize that CLEAR is more stable and robust since it can compensate for shortcomings of the base models more easily: If the aleatoric uncertainty is overestimated or underestimated, CLEAR can correct its scale by adjusting $\gamma_{1}$.
A.2 Literature that implicitly assumes $\lambda=1$
Most of the literature that combines epistemic and aleatoric uncertainty implicitly assumes that $\lambda=1$, when they simply combine epistemic and aleatoric uncertainty in the ratio 1:1. On top of the resulting uncertainty, one can use (conformal) calibration, which corresponds to CLEAR’s calibration of $\gamma_{1}$. However, in contrast to CLEAR, they do not rebalance the ratio of epistemic and aleatoric uncertainty with $\lambda$.
In practice, it is possible to underestimate the aleatoric uncertainty and overestimate epistemic uncertainty. This results in a one-dimensional calibration via $\gamma_{1}$, and consequently, narrow intervals in regions of dominating aleatoric uncertainty or in wide intervals in regions of dominating epistemic uncertainty. $\gamma_{1}$ alone cannot solve this. CLEAR can deal well with such situations by compensating for such an imbalance via $\lambda$.
@@ A.2.1 Deep Ensembles
While it is quite common to refer to deep ensembles (DE), whenever one uses an ensemble of neural networks (NNs) for uncertainty estimation, it is important to note that [22] introduced DE as a method which both estimates epistemic and aleatoric uncertainty. For regression, they train each NN with 2 outputs estimating $\mu$ and $\sigma$ via a Gaussian Maximum-Likelihood-loss (as in [34]), where $\sigma$ is responsible for the aleatoric uncertainty, which they refer to as “ambiguity in targets y for a given x”. Moreover, they use the ensemble diversity to estimate epistemic uncertainty, which they refer to as “model uncertainty”. Although DE is mainly known for its ensembling approach, [22, Table 2 in Appendix A.1] clearly shows that both the aleatoric and epistemic parts are crucial in terms of empirical performance. In their paper, the authors do not apply any calibration on top, i.e., as it is implicitly assumed that $\gamma_{2}=\gamma_{1}=1=\lambda$. However, it is common to apply (conformal) calibration on top. The commonly used calibration techniques only calibrate $\gamma_{1}$, while keeping $\lambda=1$ fixed, in contrast to CLEAR. In particular, for DE, both the diversity of the ensemble and the bias on aleatoric are very sensitive to various hyperparameters. One type of uncertainty may be strongly under-estimated. In contrast, the other type is strongly over-estimated, which motivates the need to explicitly calibrate $\lambda$ in a data-driven way.
@@
The original paper [22] is written as if applying adversarial attacks is an integral part of DEs and one of the paper’s main contributions. However, to the best of our knowledge, many practitioners refer to “DEs” without implying adversarial attacks during training, and adversarial attacks during training are seen as an optional add-on to DEs, but not an important part of DEs. Furthermore, the regression results in [22, Table 2 in Appendix A.1] do not suggest that adversarial attacks during training are particularly beneficial to the performance.
@@
Intuitively, one should estimate high epistemic uncertainty if there is a significant disagreement among the ensemble’s predictions and small epistemic uncertainty if they agree. While [40, Chapter 13] and [30] suggest using quantiles, [22] suggest using the empirical standard deviation to estimate DE’s disagreement among ensemble members’ predictions. It seems plausible that in [22]’s case of 5 ensemble members, using the standard deviation and some Gaussian assumptions can be more appropriate while in [40, Chapter 13]’s and [30]’s case of 100 or more ensemble members, quantiles can be more accurate.
@@
There exist several modifications of DE that, for example, promote the ensemble’s diversity on the function space via an additional loss term during training [46], ensemble over multiple different hyperparameters [47], or reduce the computational training cost [17, 26, 48, 49, 19, 50, 30].
@@ A.2.2 Monte Carlo Dropout
[26] originally studied Monte Carlo Dropout (MC Dropout) without explicitly modeling aleatoric uncertainty, which was then extended by [17] to also explicitly model aleatoric uncertainty. While [26] see MC Dropout as an approximation of a Bayesian neural network, one can also see it as another ensemble method. The main difference to DE is that MC Dropout only trains one NN with dropout and obtains an ensemble after training by randomly setting weights of the model to zero at inference time. Analogously to DE, MC Dropout also adds epistemic and aleatoric uncertainty in the ratio 1:1 with the same disadvantages as described before in Section A.2.1.
@@
[51] empirically concludes that ensembles have the best uncertainty and disentangling behavior of epistemic and aleatoric uncertainty. In their paper, the authors do not use any form of calibration. This would correspond to $\gamma_{2}=\gamma_{1}=1=\lambda$ in our notation. Their paper suggests a different loss function, which they call $\beta$-NLL, for training to mitigate the underestimation of aleatoric uncertainty to some extent in their experimental setting without comparing this approach to calibration. In [51, Figure 6], one can observe that even without the $\beta$-NLL, both epistemic and aleatoric uncertainty already have a good shape (that is, good relative uncertainty, see Appendix I) for deep ensembles (DE). The main problem of DE in [51, Figure 6] is that the epistemic uncertainty is too small by a very large factor (that is, poor absolute scale of uncertainty, see Appendix I), while aleatoric uncertainty has already almost the correct scaling. From our perspective, applying CLEAR on their DE would probably largely fix their problem of DE if CLEAR chooses $\gamma_{1}\approx 1$ and $\lambda\gg 1$. However, they show that for this specific experiment, $\beta$-NLL also fixes the problem. In general, where one suspects that the aleatoric or the epistemic uncertainty might be too small or too large across the domain, we strongly recommend simply applying CLEAR on top of the already trained ensemble instead of retraining all the models with a new training pipeline. The concept of CLEAR can be implemented in a few minutes and calibrates an already trained ensemble in a few seconds. We considered an interesting open problem to study if $\beta$-NLL can improve the relative epistemic and aleatoric uncertainty (see Appendix I).
@@
[14] also combines epistemic and aleatoric uncertainty in the ratio 1:1 and applies a single constant (corresponding to $\gamma_{1}$ in our notation) to scale the total predictive uncertainty, which corresponds to fixing $\lambda=1$, resulting in the same potential for problems as mentioned before.
@@ A.3.1 Gaussian Process Regression (GPR)
A.3.2 Bayesian Neural Networks (BNNs)
@@ A.3.4 TabPFN
A.4 Conformal Literature that does not account for Epistemic and Aleatoric Uncertainty